Systems Staff Meeting Notes 12/14/98

Systems Staff Meeting Notes
December 14, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Suzy Shaw

Announcements
  1. The phone in Sharon's old office is now functional. Patrick from Telecommunications fixed it last week.
  2. There has been a call to tighten up on security due to an increase in hacking & theft during the holiday season. We should be getting card swipes for the interior front & back doors, the loading dock door - & ultimately, all of our interior doors.
    Passwords - need to watch internal privileges.
    Need to make sure Outlook is completely shut down when leaving your work station.
    Doors need to stay locked.
    Sign-ons & permissions for NT servers - Suzy is reviewing, access will be changing - may cause lock-outs. Will try to expand user authority to grant permissions. Current NT doesn't enforce disk limitations - path statement at autoexec level has been a problem - sort of distributed administration.
    Winston proposes that we do not use general sign-ons on NT's. This would require changes to services - will need to take the time to investigate & make the changes.
  3. Tom Cetwinski recently announced his departure. He will be leaving soon, a search committee will be formed to find his replacement.
  4. An offer will soon be made to a selected individual for the Programmer position.
  5. Ashley has a training session on hiring scheduled with Erika in Personnel, then he will move on the filling of the Repair Tech position.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Replacement hard drive disks have not yet been received.
  2. We need to devise a signout procedure for hardware/software tracking.
Networking
  1. Criser POP - unsure if it is complete, we have not been moved yet.
  2. There was another network intrusion in a branch library. Filters prevented the action from success. Bill will be called next time it occurs.
  3. New mail server - maybe this weekend? The biggest question we now face is how to switch the listserves over to NT7. May need to change DNS pointers.
  4. We will be upgrading substandard wiring in room 219 West, Resource Services North tower & MSL closet. In the MSL closet, we'll need to change from 400' jumpers - will need to do the patch panels to switch. Will also need tools & labeling, documentation - may require some down time. Cat 5 interconnects, ends & racks need to be installed. May get to MSL right after Christmas.
  5. The Control Unit in MSL is dripping something - need to call IBM service for it, as the substance could be carcinogenic.
  6. Ashley found a KillConn command for the Smathers Novell server to kill the 'not logged-in' spaces. It can be scheduled to run hourly.
Software Projects Status
  1. Mellon Demo - Erich, Bill, Martha & Winston met - some cosmetic changes were identified, nothing major.
  2. FCLA backup ran.
  3. All of the Macs have been removed from Personnel & Sharon's old office. Mac files moved to NT3 okay.
  4. OPS Tracker - Resource Services began inputting around Thanksgiving time, but then stopped.
  5. Newspaper Project - LaVerne has the prototype, will migrate from FoxPro to Access when she's ready.
  6. ATC Standalone Catalog - needs to be started early in the next calendar year - could do it on the server.
  7. ILL Borrowing - will have major Y2K problems if the conversion is not complete. CLEO meeting is being held today to discuss parallel system for ILL.
DP Coordinator
  1. Students will dwindle down to nothing by the end of the week.
  2. Compustat tapes - awaiting detailed email from Sharon Wheeler at NERDC describing what has been done to the tapes so that Bill can contact Compustat.
  3. Spanish proverbs disks which were compiled by a Romance Language professor have been delivered to Bill. Debra will make 2 backups and Bill will have the task of trying to decipher the information.
  4. SPOT tapes - Scott Smith had an email exchange with a salesman who authorized the backup procedures in writing. Geography's writeable CD project is on go ahead status. The media will be ordered.
Calendar
  1. Rich Bennett will be out this afternoon.
  2. Library Christmas Party will be Wednesday, Dec. 16.
  3. Systems Staff Christmas Luncheon will be Friday, December 18.
  4. Ashley out December 21 & 22.
  5. Rich out December 21-28, back December 29.
  6. Debra out December 22-28, back December 29.
  7. Bill out December 21-31, back January 4.
  8. Bobbie out December 24.
  9. Winston out December 29.


Back to the Systems Homepage

Last updated February 24, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu